# Electroweak Phase Transition in Two Scalar Singlet Model with pNGB Dark Matter

### v1

@@ VII Conclusion
In this work, we explored the dark matter (DM) phenomenology and the dynamics of strong first-order phase transitions (SFOPT) within a model featuring pseudo-Nambu-Goldstone boson (pNGB) dark matter, realized through a $\mathcal{Z_{\rm 3}}$-symmetric complex scalar, alongside a $\mathcal{Z_{\rm 2}}$-symmetric real scalar both of which are singlet under the $SU(2)_{L}$ symmetry. Our analysis revealed two primary patterns of strong first-order phase transitions, including transitions along both SM+BSM Higgs direction and BSM Higgs direction only. However, stringent DM direct detection constraints significantly constrain the parameter space, excluding many points where SFOPT was observed. This exclusion underscores the challenge of achieving SFOPT within a viable parameter space, where the complex scalar DM is frequently under-abundant. The dynamics of phase transitions and their associated gravitational wave signatures within a theoretical framework featuring a $\mathcal{Z}_{3}$-symmetric scalar field have been investigated. By introducing a $\mathcal{Z}_{2}$ symmetric real scalar field into this framework, we have identified new parameters— $m_{H_{3}},v_{\phi}$ and $\sin\alpha_{2}$ —that significantly affect the strength of the phase transition along the Standard Model (SM) Higgs direction.
To explore the implications of these additional parameters, we analyzed seven benchmark scenarios. Three of these scenarios exhibit strong first-order phase transitions (SFOPT) along both the SM and the BSM Higgs field directions, while the remaining four demonstrate SFOPT exclusively along the BSM Higgs direction. The benchmark Point BP1 shows a single-step first-order phase transition with pronounced transitions along all three field directions: $H_{1},H_{2}$ and $H_{3}$. The benchmark Points BP2 depicts a two-step first-order phase transitions: the initial step is characterized by a strong transition along the $H_{2}$ direction, while the subsequent step involves strong transitions along the $H_{1}$ field directions. On the other BP3 also exhibits 2 step FOPT similar to BP2, with strong transitions along the $H_{1}$ (first step) and $H_{2}$ field direction in the second step. The benchmark Points BP4, BP5, and BP6 all showcase a three-step phase transition, with a strong first-order transition occurring exclusively along the $H_{2}$ direction in the initial step. Finally, the benchmark Point BP7 demonstrates a two-step phase transition, characterized by a strong first-order transition along the $H_{2}$ direction in the first step.
@@
Through our analysis, we have shown that part of the model parameters that produce strong first order phase transitions are in direct conflict with the dark matter direct detection constraints provided by the LZ collaboration. For example, the LZ observation rules out the benchmark points BP2, BP3, and BP7, though these benchmark points are consistent with the Xenon 1T limit. The final four benchmark points, that exhibit SFOPT along the SM Higgs field direction show highly underabundant relic density. Even when SFOPT occurs along the Beyond Standard Model (BSM) Higgs field direction, the resulting scalar DM abundance often falls well below the $2\sigma$ range of the observed relic density. Notably, the direct detection cross-section does not vanish at tree level for zero momentum transfer due to the presence of a single symmetry-breaking scale parameter $\mu_{3}$. This parameter is also critical in establishing a barrier in the tree-level potential, enabling first-order phase transitions (FOPT) in the singlet Higgs field direction within a substantial portion of the parameter space while still evading DM direct detection constraints.
Finally, we have checked the testability of the final benchmark points which satisfy both dark matter direct and indirect detection constraints by evaluating the energy density of the gravitational wave spectrum corresponding to those BPs. We have found that the highest peak amplitude of the GW corresponds to BP4, followed by BP5 and BP6 respectively. While the gravitational wave spectrum generated by SFOPT along the SM Higgs direction (BP1) is unlikely to be detectable in near-future experiments, the first-order phase transition (FOPT) signatures along the BSM Higgs direction present a promising avenue for detection in upcoming gravitational wave observatories like DECIGO correlation, U-DECIGO and U-DECIGO correlation.

### v2

@@ VII Conclusion
In this work, we explored the dark matter (DM) phenomenology and the dynamics of strong first-order phase transitions (SFOPT) within a model featuring pseudo-Nambu-Goldstone boson (pNGB) dark matter, realized through a $\mathcal{Z_{\rm 3}}$-symmetric complex scalar, alongside a $\mathcal{Z_{\rm 2}}$-symmetric real scalar both of which are singlet under the $SU(2)_{L}$ symmetry. Our analysis revealed two primary patterns of strong first-order phase transitions, including transitions along both SM+BSM Higgs direction and BSM Higgs direction only. However, stringent DM direct detection constraints significantly constrain the parameter space, excluding many points where SFOPT was observed. This exclusion underscores the challenge of achieving SFOPT within a viable parameter space, where the complex scalar DM is frequently under-abundant. The dynamics of phase transitions and their associated gravitational wave signatures within a theoretical framework featuring a $\mathcal{Z}_{3}$-symmetric scalar field have been investigated. By introducing a $\mathcal{Z}_{2}$ symmetric real scalar field into this framework, we have identified new parameters— $m_{H_{3}},v_{\phi}$ and $\sin\alpha_{2}$ —that significantly affect the strength of the phase transition along the SM Higgs direction.
To explore the implications of these additional parameters, we analyzed seven benchmark scenarios. Three of these scenarios exhibit SFOPT along both the SM and the BSM Higgs field directions, while the remaining four demonstrate SFOPT exclusively along the BSM Higgs direction. The benchmark Point BP1 shows a single-step first-order phase transition with pronounced transitions along all three field directions: $H_{1},H_{2}$ and $H_{3}$. The benchmark Points BP2 depicts a two-step first-order phase transitions: the initial step is characterized by a strong transition along the $H_{2}$ direction, while the subsequent step involves strong transitions along the $H_{1}$ field directions. On the other BP3 also exhibits 2 step FOPT similar to BP2, with strong transitions along the $H_{1}$ (first step) and $H_{2}$ field direction in the second step. The benchmark Points BP4, BP5, and BP6 all showcase a three-step phase transition, with a strong first-order transition occurring exclusively along the $H_{2}$ direction in the initial step. Finally, the benchmark Point BP7 demonstrates a two-step phase transition, characterized by a strong first-order transition along the $H_{2}$ direction in the first step.
@@
Through our analysis, we have shown that part of the model parameters that produce strong first order phase transitions are in direct conflict with the dark matter direct detection constraints provided by the LZ collaboration. For example, the LZ observation rules out the benchmark points BP2, BP3, and BP7, though these benchmark points are consistent with the Xenon 1T limit. The final four benchmark points, that exhibit SFOPT along the SM Higgs field direction show highly underabundant relic density. Even when SFOPT occurs along the BSM Higgs field direction, the resulting scalar DM abundance often falls well below the $2\sigma$ range of the observed relic density. Notably, the direct detection cross-section does not vanish at tree level for zero momentum transfer due to the presence of a single symmetry-breaking scale parameter $\mu_{3}$. This parameter is also critical in establishing a barrier in the tree-level potential, enabling first-order phase transitions in the singlet Higgs field direction within a substantial portion of the parameter space while still evading DM direct detection constraints.
Finally, we have checked the testability of the final benchmark points which satisfy both dark matter direct and indirect detection constraints by evaluating the energy density of the gravitational wave spectrum corresponding to those BPs. We have found that the highest peak amplitude of the GW corresponds to BP4, followed by BP5 and BP6 respectively. While the gravitational wave spectrum generated by SFOPT along the SM Higgs direction (BP1) is unlikely to be detectable in near-future experiments, the first-order phase transition signatures along the BSM Higgs direction present a promising avenue for detection in upcoming gravitational wave observatories like DECIGO correlation, U-DECIGO and U-DECIGO correlation.